10g of Sodium hydroxide reacts with 6g of Hydrochloric acid to form 12g of Sodium Chloride along with some water. Calculate the mass of water formed in the reaction according to the given information. State the law used by you in calculating this mass.
1
Expert's answer
2020-12-23T04:05:51-0500

Q152342

 

10g of Sodium hydroxide reacts with 6g of Hydrochloric acid to form 12g of Sodium Chloride along with some water. Calculate the mass of water formed in the reaction according to the given information. State the law used by you in calculating this mass.


Solution:


Convert 10g of NaOH and 6 g of HCl to moles by using their molar mass.


molar mass of NaOH = 1 * atomic mass of Na + 1 * atomic mass of O + 1 * atomic mass of H

= 1 * 22.99 g/mol + 1 * 15.999 g/mol + 1 * 1.0079 g/mol

= 39.9969g/mol



moles of NaOH = "=10g NaOH * \frac{1mol NOH}{39.9969g NaOH} = 0.2500 mol NaOH" ;


molar mass of HCl = 1 * atomic mass of H + 1 * atomic mass of Cl

= 1 * 1.0079g/mol + 1 * 35.453g/mol

= 36.4609 g/mol


moles of HCl = "=6g HCl * \frac{1molHCl}{36.4609g HCl} = 0.1646 mol NaOH ;"


The reaction of NaOH(aq) and HCl (aq) is given as


NaOH (aq) + HCl (aq) -----> NaCl (aq) + H2O(l) ;


We can see that the mol to mol ratio of NaOH and HCl in the given reaction is 1 :1.


So same moles of NaOH and HCl combine to give NaCl and H2O.


0.1646 mol HCl will react with 0.1646 mol of NaOH. The remaining NaOH will remain unreacted

in the solution.


1) To find moles of H2O formed


The moles of product formed depends on the moles of Limiting reactant present. We now that

HCl is a limiting reactant here and NaOH is present is in excess.

So moles of H2O depend on moles of HCl present.


The mol to mol ratio of HCl and H2O is 1: 1.


so moles of H2O formed = "= 0.1646 mol HCl *" "\frac{1mol H_2O}{1mol HClg } = 0.1646 mol H_2O;"



2) To find grams of H2O formed


Convert 0.1646 mol H2O to grams using the molar mass of H2O.


molar mass of H2O = 2 * atomic mass of H + 1 * atomic mass of O

= 2 * 1.0079 g/mol + 1 * 15.999g/mol

= 2.0159g/mol + 15.999 g/mol

= 18.0148 g/mol


grams of H2O = 0.1646 mol H2O * "\frac{18.0148g H_2O}{1mol H_2O} = 2.965 g H_2O ;"



In the given question, the quantity with the least number of the significant figure is 6 g HCl. It is in 1 significant figure, so we round up our answer to 1 significant figure.


Hence the mass of H2O formed is 3 grams.


We have used the Law of Conservation of mass in calculating the mass.


Law of Conservation of mass: For a chemical process in a closed system, the mass of reactants must equal the mass of products formed.


We used the mole ratio of limiting reactant HCl and H2O for finding the moles of H2O formed

and then used the molar mass of H2O for converting 'moles of H2O' to ' grams of H2O'.